### v1

@@ 1 Introduction
However, the research efforts of FIDL have gradually split into four relatively independent research domains over time. 1) Deepfake detection (Deepfake) [37, 26, 25, 8, 38, 4, 66, 42, 28, 33, 54, 71, 49, 65, 67]: detects manipulations such as face swapping, expression editing, or feature replacement. 2) Image manipulation detection/localization (IMDL) [5, 17, 29, 35, 70, 53]: detects and localizes the tampering in natural images. 3) AI-Generated Image Detection (AIGC) [40, 18, 63, 59]: detects the images fully generated by deep generative models such as Stable Diffusion [46]. 4) Document Image Manipulation Localization (Document)  [48, 43, 6, 12, 52]: localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
Although these domains have become isolated due to differences in application scenarios, manipulation types, and detection methods, there are still overlaps and similarities among them. As vision tasks, these four domains almost universally adopt SoTA detection or segmentatoin models as pre-trained backbones. Further, since the creators of fake images typically aim to preserve semantically plausible and realistic content, all four domains have placed considerable emphasis on designing low-level visual feature extractors to capture subtle, non-semantic discrepancies for reliable detection. Some research methodologies, such as contrastive learning, are commonly employed across these areas to mine discriminative features. We summarize SoTAs in four domains of the backbone, artifacts strategy, output type, and contribution in Table 1. The differences cause the four FIDL domains to become fragmented, but the similarities call for a unified perspective to understand them cohesively.
@@
Deepfake Capsule-Net [37](ICASSP19) VGG [50] Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
@@
RECCE [4](CVPR22) Xception [7] Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
@@
SPSL [28](CVPR21) Xception [7] Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
@@
UCF [66](ICCV23) Xception [7] Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
@@
SBI [49](CVPR22) EfficientNet [55] Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
@@
IMDL MVSS-Net [62](ICCV21) Resnet [19] BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
@@
CAT-Net [22](IJCV22) HRNet [56] DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
@@
PSCC-Net [29](TCSVT22) HRNet [56] Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
@@
Trufor [17](CVPR23) Seformer [64] High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
@@
IML-ViT [35](Arxiv) ViT [13] BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
@@
Mesorch [70](AAAI25) Conv. [31],Segfor. [64] DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
@@
AIGC Dire [59](ICCV23) Resnet [19] Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
@@
DualNet [63](APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
@@
HiFiNet [18](CVPR23) HRNet [56] Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
@@
Synthbuster [2](OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
@@
UnivFD [40](CVPR23) CLIP-ViT [45] None Label Use pretrained vision-language model features for unified detection.
@@
Document CAFTB [52](TOMM24) Resnet [19] SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
@@
TIFDM [12](TCE24) Resnet [19] None Mask Proposes a robust network with multiscale attention.
@@
DTD [43](CVPR23) Conv. [31], Swin.[30] Frequency Mask Proposes DTD with frequency head and multi-view decoder.
@@
FFDN [6](ECCV24) ConvNext [31] Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
@@
Although individual benchmarks exist in some domains, such as DeepfakeBench [67] for Deepfake and IMDLBenCo [36] for IMDL, a unified benchmark for all domains in FIDL remains blank. The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability. Domain silos lead to redundant and uneven research across existing FIDL fields, and difficulty in establishing a general and unified FIDL approach, severely hindering the development of the entire FIDL field.
@@
To this end, we propose ForensicHub, which: 1) proposes a modular and configuration-driven architecture that decomposes forensic pipelines into interchangeable components across datasets, transforms, models, and evaluators, allowing flexible composition across all domains; 2) fully implements 10 baseline models (3 of which are reproduced from scratch), 6 backbones, 2 new benchmarks for AIGC and Doc, and integrates 2 existing benchmarks of DeepfakeBench and IMDLBenCo through an adapter-based design. With the above efforts, ForensicHub serves as the first unified benchmark and codebase for all-domain fake image detection and localization. Building on ForensicHub, we establish an image forensic fusion protocol (IFF-Protocol) evaluation mechanism that supports unified training and testing of diverse forensic models across tasks. We conduct deep analysis on 8 issues that are of particular interest to researchers but have not yet been thoroughly investigated, offering new insights into FIDL model architecture, dataset characteristics, and evaluation standards. The introduction of ForensicHub bridges all domains within FIDL, breaks down domain silos, and inspires future breakthroughs.
@@ 2 Related Works
Fake image detection and localization span four sub-tasks, summarized in Appendix C. Despite rapid progress, a unified benchmark is lacking—each task uses isolated pipelines, limiting cross-task comparison.
DeepfakeBench [67] is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models. IMDLBenCo [36] is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol. AIGCDetectBenchmark [68] is a repository for experiments on 9 AI-generated image detection methods.
@@ 3 ForensicHub
To accommodate different forensic tasks, ForensicHub is designed as a modular architecture consisting of four main components: Datasets, Transforms, Models, and Evaluators. Datasets handle the data loading process and are required to return fields that conform to the ForensicHub specification. Transforms handle the data pre-processing and augmentation for different tasks. Models, through alignment with Datasets and unified output, allow for the inclusion of various state-of-the-art image forensic models. Evaluators cover commonly used image- and pixel-level metrics for different tasks, and are implemented with GPU acceleration to improve evaluation efficiency during training and testing.
@@
To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design [14] that ensures seamless integration with DeepfakeBench [67] and IMDLBenCo [36], two widely used benchmarks. This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol. This unified infrastructure simplifies cross-task benchmarking, supports reproducibility, and enables consistent evaluation across domains.
@@
Datasets used in this paper are: Faceforencs++ [47], Celeb-DF [27], DFD [9], FaceShifter [23] and UADFV [26] for Deepfake, CASIA [11], COVERAGE [60], Columbia [20], IMD2020 [39], NIST16 [16], CocoGlide [17], and Autosplice [21] for IMDL, DiffusionForensics [59], GenImage [69] for AIGC, Doctamper [43], T-SROIE [58], OSTF [44], TPIC-13 [57], RTM [32] for Doc. A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
@@
Deepfake FaceForensics++ [47] 2019 Face 45,388 127,209 Label Automatic Train,Test
Celeb-DF-v1 [27] 2020 Face 1,203 1,933 Label Automatic Train,Test
Celeb-DF-v2 [27] 2020 Face 5,620 10,800 Label Automatic Train,Test
DeepFakeDetection [15] 2019 Face 10,741 91,800 Label Automatic Train,Test
DFDC [9] 2020 Face 63,265 68,851 Label Automatic Train,Test
DFDCP [10] 2019 Face 5,901 11,321 Label Automatic Train,Test
FaceShifter [24] 2020 Face 4,479 4,479 Label Automatic Train,Test
UADFV [26] 2018 Face 1,548 1,551 Label Automatic Train,Test
IMDL CASIAv1 [11] 2013 General 0 920 Label,Mask Manual Train
CASIAv2 [11] 2013 General 7,491 5,123 Label,Mask Manual Train
COVERAGE [60] 2016 General 100 100 Label,Mask Manual Train
Columbia [20] 2006 General 183 180 Label,Mask Automatic Train
IMD2020 [11] 2020 General 414 2,010 Label,Mask Manual Train
NIST16 [16] 2019 General 873 564 Label,Mask Manual Train
CocoGlide [17] 2023 General 512 512 Label,Mask Automatic Train
Autosplice [21] 2023 General 2,273 3,621 Label,Mask Manual Train
AIGC DiffusionForensics [59] 2023 General 134,000 481,200 Label Automatic Train,Val,Test
GenImage [69] 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
Document Doctamper [43] 2023 Document 0 170,000 Mask Automatic Train,Test
OSTF [44] 2025 Document 16,405 3,685 Mask Manual Train,Test
RealTextManipulation [32] 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
T-SROIE [58] 2022 Document 21,268 4,326 Mask Manual Train,Test
Tampered-IC13 [57] 2022 Document 2,810 1,228 Mask Manual Train,Test
Models used in this paper are: Capsule-Net [37], RECCE [4], SPSL [28], UCF [66], and SBI [49] for Deepfake detection, MVSS-Net [5], CAT-Net [22], PSCC-Net [29], Trufor [17], IML-ViT [35], and Mesorch [70] for image manipulation and localization, Dire [59], DualNet [63], HiFiNet [18], Synthbuster [2], and UnivFD [40] for AIGC detection, DTD [43], FFDN [6], CAFTB [52], TIFDM [12] for document detection. These methods are from official repositories and our reimplementations. In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet [19], Xception [7], EfficientNet [55], Segformer [64], Swin Transformer [30], and ConvNext [31]. Details about models can be found in Appendix D.2.
@@ 4 Benchmarks
In addition to being fully compatible with existing benchmarks, DeepfakeBench [67] and IMDLBenCo [36], ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases. We propose two protocols for two domains to evaluate generalization.

### v2

@@ 1 Introduction
However, the research efforts of FIDL have gradually split into four relatively independent research domains over time. 1) Deepfake detection (Deepfake) nguyen2019capsule ; li2018uadfv ; li2020xray ; dang2020ffd ; ni2022core ; cao2022recce ; yan2023ucf ; han2023fcd ; qian2020f3net ; liu2021spsl ; luo2021srm ; sun2022SIA ; zhuang2022uia ; shiohara2022sbi ; yan2024LSDA ; yan2023deepfakebench : detects human-centric manipulations such as face swapping, expression editing, or feature replacement. 2) Image manipulation detection/localization (IMDL) liu2020d ; MVSS_2021 ; trufor2023 ; liu2022pscc ; NCL_IML_2023 ; ma2023iml ; yu2024diffforensics ; zhu2025mesoscopic ; sparseViT_2025 : detects and localizes the tampering in natural images. 3) AI-Generated Image Detection (AIGC) univfd ; HiFi-Net2023 ; dualnet ; wang2023dire ; zhang2025towards : detects the images fully generated by deep generative models such as Stable Diffusion stablediffusion . 4) Document Image Manipulation Localization (Document)  psnet ; qu2023doctamper ; chen2024ffdn ; dong2024tifdm ; song2025caftb ; yu2023learning ; li2026document : localizes the forgery of various forms of document images, including receipts, certificates, and identification materials, with a particular focus on detecting modifications to the printed text.
Although these domains have become isolated due to differences in application scenarios, manipulation types, and detection methods, there are still overlaps and similarities among them. As vision tasks, these four domains almost universally adopt SoTA detection or segmentation models as pre-trained backbones. Further, since the creators of fake images typically aim to preserve semantically plausible and realistic content, all four domains have placed considerable emphasis on designing low-level visual feature extractors to capture subtle, non-semantic discrepancies for reliable detection. Some research methodologies, such as contrastive learning, are commonly employed across these areas to mine discriminative features.
We summarize SoTAs in four domains of the backbone, artifacts strategy, output type, and contribution in Table 1. The differences cause the four FIDL domains to become fragmented, but the similarities call for a unified perspective to understand them cohesively.
@@
Deepfake Capsule-Net nguyen2019capsule (ICASSP19) VGG VGG_2015 Dynamic Routing Label Proposes a capsule network with dynamic routing and a VGG19 backbone.
@@
RECCE cao2022recce (CVPR22) Xception chollet2017xception Reconstruction Label Proposes a graph-based framework leveraging reconstruction differences
@@
SPSL liu2021spsl (CVPR21) Xception chollet2017xception Phase Spectrum Label Proposes phase-spectrum fusion with Xception for face forgery detection.
@@
UCF yan2023ucf (ICCV23) Xception chollet2017xception Multi-task Disentanglement Label Proposes multi-task disentanglement with Xception for deepfake generalization.
@@
SBI shiohara2022sbi (CVPR22) EfficientNet tan2019efficientnet Frequency,Blending Boundaries Label Proposes self-blended images to improve deepfake detection generalization.
@@
IMDL MVSS-Net Mantra_2019 (ICCV21) Resnet Resnet_2016 BayarConv,Sobel Label,Mask Exploit noise and boundary artifacts via multi-view learning for manipulation detection.
@@
CAT-Net CAT-Net2022 (IJCV22) HRNet HRNet DCT Mask Fuse RGB and DCT streams to learn compression artifacts for splice localization.
@@
PSCC-Net liu2022pscc (TCSVT22) HRNet HRNet Multi-Resolution Conv Label,Mask Progressively refine masks with spatio-channel correlations for high-reso localization.
@@
Trufor trufor2023 (CVPR23) Seformer SegFormer_2021 High-Reso,Multi-scale,Edge Label,Mask Fuse RGB and learned noise fingerprints to detect manipulations as anomalies.
@@
IML-ViT ma2023iml (Arxiv) ViT ViT_2021 BayarConv,SRM Filter Mask Use ViT with high-reso, multi-scale edge-aware design for manipulation localization.
@@
Mesorch zhu2025mesoscopic (AAAI25) Conv. Convnet_2022 ,Segfor. SegFormer_2021 DCT Mask Fuse micro and macro cues for mesoscopic image manipulation localization.
@@
AIGC Dire wang2023dire (ICCV23) Resnet Resnet_2016 Diffusion Reconstruction Label Use reconstruction error of diffusion for diffusion-generated images detection.
@@
DualNet dualnet (APSIPA23) CNN SRM,Low Frequency Label Fuse SRM residual and low-frequency content streams for AIGC detection.
@@
HiFiNet HiFi-Net2023 (CVPR23) HRNet HRNet Multi-branch Feature Extractor Label,Mask Learn hierarchical fine-grained representations of forgery attributes.
@@
Synthbuster bammey2023synthbuster (OJSP23) None Fourier Transform Label Leverage spectral artifacts in the frequency domain for diffusion detection.
@@
UnivFD univfd (CVPR23) CLIP-ViT clip None Label Use pretrained vision-language model features for unified detection.
@@
Document CAFTB song2025caftb (TOMM24) Resnet Resnet_2016 SRM Mask Proposes CAFTB-Net with dual-branch and cross-attention.
@@
TIFDM dong2024tifdm (TCE24) Resnet Resnet_2016 None Mask Proposes a robust network with multiscale attention.
@@
DTD qu2023doctamper (CVPR23) Conv. Convnet_2022 , Swin.Swin_2021 Frequency Mask Proposes DTD with frequency head and multi-view decoder.
@@
FFDN chen2024ffdn (ECCV24) ConvNext Convnet_2022 Wavelet, Frequency Mask Proposes FFDN combining visual enhancement and frequency decomposition
@@
Although individual benchmarks exist in some domains, such as DeepfakeBench yan2023deepfakebench for Deepfake and IMDLBenCo ma2024imdl for IMDL, a unified benchmark for all domains in FIDL remains blank. The absence of such a unified benchmark results in significant domain silos, where each domain independently constructs its datasets, models, and evaluation protocols without interoperability. Domain silos lead to redundant and uneven research across existing FIDL fields, and difficulty in establishing a general and unified FIDL approach, severely hindering the development of the entire FIDL field.
Besides, in real-world scenarios, it is often impossible to predetermine the type of manipulation (deepfake, imdl, aigc and document) present in an image, making unified detection particularly important for users.
@@
To this end, we propose ForensicHub, which: 1) proposes a modular and configuration-driven architecture that decomposes forensic pipelines into interchangeable components across datasets, transforms, models, and evaluators, allowing flexible composition across all domains; 2) fully implements 10 baseline models (3 of which are reproduced from scratch), 6 backbones, 2 new benchmarks for AIGC and Doc, and integrates 2 existing benchmarks of DeepfakeBench and IMDLBenCo through an adapter-based design.
With the above efforts, ForensicHub serves as the first unified benchmark and codebase for all-domain fake image detection and localization. Building on ForensicHub, we establish an image forensic fusion protocol (IFF-Protocol) evaluation mechanism that supports unified training and testing of diverse forensic models across tasks. We conduct deep analysis on 8 issues that are of particular interest to researchers but have not yet been thoroughly investigated, offering new insights into FIDL model architecture, dataset characteristics, and evaluation standards. The introduction of ForensicHub bridges all domains within FIDL, breaks down domain silos, and inspires future breakthroughs.
@@ 2 Related Works
Fake image detection and localization encompass four sub-tasks: 1) Deepfake Detection, 2) Image Manipulation Detection and Localization, 3) AI-Generated Image Detection, and 4) Document Image Manipulation Localization. The characteristics of each task are summarized in Appendix C. Despite rapid progress, a unified benchmark is lacking—each task uses isolated pipelines, limiting cross-task comparison.
Despite the rapid development of these tasks, there is a lack of a unified benchmark, with some task having its isolated benchmark, creating barriers between them.
DeepfakeBench yan2023deepfakebench is a Deepfake detection benchmark specifically designed to address the lack of uniformity in data processing pipelines, leading to inconsistent data inputs for detection models. IMDLBenCo ma2024imdl is a benchmark and codebase for IMDL, aiming to compare IMDL models through a unified training and evaluation protocol. AIGCDetectBenchmark EkkoAIGCDetectBenchmark is a repository for experiments on 9 AI-generated image detection methods.
@@ 3 ForensicHub
To accommodate different forensic tasks, ForensicHub is designed as a modular architecture consisting of four main components: Datasets, Transforms, Models, and Evaluators. 1) Datasets handle the data loading process and are required to return fields that conform to the ForensicHub specification. 2) Transforms handle the data pre-processing and augmentation for different tasks. 3) Models, through alignment with Datasets and unified output, allow for the inclusion of various state-of-the-art image forensic models. 4) Evaluators cover commonly used image- and pixel-level metrics for different tasks, and are implemented with GPU acceleration to improve evaluation efficiency during training and testing.
@@
To enable broad interoperability and reduce duplication of effort, ForensicHub adopts an adapter-based design gamma1995design that ensures seamless integration with DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , two widely used benchmarks. This mechanism allows users to reuse existing models and datasets without major modification, while also supporting the definition of new models and benchmarks within ForensicHub under the unified protocol. This unified infrastructure simplifies cross-task benchmarking, supports reproducibility, and enables consistent evaluation across domains.
@@
Datasets used in this paper are: FaceForensics++ rossler2019faceforensics++ , Celeb-DF li2020celeb , DFD dolhansky2020dfd , FaceShifter li2019faceshifter and UADFV li2018uadfv for Deepfake; CASIA CASIA_2013 , COVERAGE Coverage_2016 , Columbia Columbia_2006 , IMD2020 IMD20_2020 , NIST16 NIST16_2019 , CocoGlide trufor2023 , and Autosplice jia2023autosplice for IMDL; DiffusionForensics wang2023dire , GenImage zhu2023genimage for AIGC; Doctamper qu2023doctamper , T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , RTM luo2025rtm for Doc. A brief summary of each dataset is provided in Table 2, with more details in Appendix D.1.
@@
Deepfake FaceForensics++ rossler2019faceforensics++ 2019 Face 45,388 127,209 Label Automatic Train,Test
Celeb-DF-v1 li2020celeb 2020 Face 1,203 1,933 Label Automatic Train,Test
Celeb-DF-v2 li2020celeb 2020 Face 5,620 10,800 Label Automatic Train,Test
DeepFakeDetection dfd2019 2019 Face 10,741 91,800 Label Automatic Train,Test
DFDC dolhansky2020dfd 2020 Face 63,265 68,851 Label Automatic Train,Test
DFDCP dolhansky2019deepfake 2019 Face 5,901 11,321 Label Automatic Train,Test
FaceShifter li2020advancing 2020 Face 4,479 4,479 Label Automatic Train,Test
UADFV li2018uadfv 2018 Face 1,548 1,551 Label Automatic Train,Test
IMDL CASIAv1 CASIA_2013 2013 General 0 920 Label,Mask Manual Train
CASIAv2 CASIA_2013 2013 General 7,491 5,123 Label,Mask Manual Train
COVERAGE Coverage_2016 2016 General 100 100 Label,Mask Manual Train
Columbia Columbia_2006 2006 General 183 180 Label,Mask Automatic Train
IMD2020 CASIA_2013 2020 General 414 2,010 Label,Mask Manual Train
NIST16 NIST16_2019 2019 General 873 564 Label,Mask Manual Train
CocoGlide trufor2023 2023 General 512 512 Label,Mask Automatic Train
Autosplice jia2023autosplice 2023 General 2,273 3,621 Label,Mask Manual Train
AIGC DiffusionForensics wang2023dire 2023 General 134,000 481,200 Label Automatic Train,Val,Test
GenImage zhu2023genimage 2023 General 1,331,167 1,350,000 Label Automatic Train,Test
Document Doctamper qu2023doctamper 2023 Document 0 170,000 Mask Automatic Train,Test
OSTF qu2025ostf 2025 Document 16,405 3,685 Mask Manual Train,Test
RealTextManipulation luo2025rtm 2025 Document 52,314 17,423 Mask Automatic+Manual Train,Test
T-SROIE wang2022tsroie 2022 Document 21,268 4,326 Mask Manual Train,Test
Tampered-IC13 wang2022tpic 2022 Document 2,810 1,228 Mask Manual Train,Test
Models used in this paper are: Capsule-Net nguyen2019capsule , RECCE cao2022recce , SPSL liu2021spsl , UCF yan2023ucf , and SBI shiohara2022sbi for Deepfake detection; MVSS-Net MVSS_2021 , CAT-Net CAT-Net2022 , PSCC-Net liu2022pscc , Trufor trufor2023 , IML-ViT ma2023iml , and Mesorch zhu2025mesoscopic for image manipulation and localization; Dire wang2023dire , DualNet dualnet , HiFiNet HiFi-Net2023 , Synthbuster bammey2023synthbuster , and UnivFD univfd for AIGC detection; DTD qu2023doctamper , FFDN chen2024ffdn , CAFTB song2025caftb , TIFDM dong2024tifdm for document detection. These methods are from official repositories and our reimplementations. In addition, ForensicHub also selects 6 commonly used backbones in visual tasks, which are: Resnet Resnet_2016 , Xception chollet2017xception , EfficientNet tan2019efficientnet , Segformer SegFormer_2021 , Swin Transformer Swin_2021 , and ConvNext Convnet_2022 . Details about models can be found in Appendix D.2.
@@ 4 Benchmarks
In addition to being fully compatible with existing benchmarks, DeepfakeBench yan2023deepfakebench and IMDLBenCo ma2024imdl , ForensicHub further extends standardization efforts by introducing unified evaluation protocols for the AIGC and Document domains—two areas that previously lacked widely accepted benchmarks and codebases. We propose two protocols for two domains to evaluate generalization.